OpenStudy (anonymous):

Can can someone lease check my answers and tell me if im right??? You have four fixed-volume containers at STP. Container A has 0.5 mol of gas in 11.2 L. Container B has 2 mol of gas in 22.4 L. Container C has 1 mol of gas in 22.4 L. Container D has 2 mol of gas in 11.2 L. Which of the four containers have equal pressures? (Points : 3) A and C A and B B and C C and D *****MY answer is A

OpenStudy (anonymous):

Why does opening the air valve of a tire at a constant temperature decrease the pressure? (Points : 3) The temperature decreases. The volume decreases. The atmospheric pressure decreases. The number of molecules decreases. ****My answer is D

OpenStudy (anonymous):

Which gas law has the assumptions of a fixed amount and volume of a gas? (Points : 3) Avogadro's law Boyle's law Charles's law Gay-Lussac's law ***** Im not sure about this one I think its C

OpenStudy (anonymous):

Which of the following is the mathematical representation of Gay-Lussac's law? (Points : 3) P1/T1 = P2/T2 P1V1 = V2/T2 V1/T1 = V2/T2 P1V1 = P2V2 ****i dont know

OpenStudy (anonymous):

A student increases the temperature of a 200 cm3 balloon from 60°C to 120°C. What will the new volume of the balloon be? (Be careful with units.) (Points : 3) 100 cm3 236 cm3 200 cm3 400 cm3 My answer is B

OpenStudy (anonymous):

In a reaction between vinegar and antacid tablets, the antacid is the limiting reagent. At constant pressure and temperature, three tablets produce 600 cm3 of gas. What volume will four tablets produce? (Points : 3) 300 cm3 600 cm3 800 cm3 1,200 cm3 ***my answer is C

OpenStudy (anonymous):

You breathe in 12.0 L of pure oxygen at 298 K and 1,000 kPa to fill your lungs. How many moles of oxygen did you take in? (Use the ideal gas law: PV = nRT where R = 8.31 L-kPa/mol-K.) (Points : 3) 0.05 moles 1.21 moles 2.42 moles 4.84 moles ***my answer is D

OpenStudy (anonymous):

Consider a gas at STP in a container of 22.4 L. If you apply the ideal gas law, what is the approximate value of n? (Points : 3) 224 1 8.31 0.5 ****im not sure

OpenStudy (anonymous):

Water boils at 100°C. What is that temperature on the Kelvin scale? (Points : 3) 213 K 325 K 333 K 373 K ***my answer is D

OpenStudy (aaronq):

too many q's